Absorbance [CrO₄²⁻] [Ag⁺] Ksp of Ag2CrO4
0.214 1.91 x 10-4 3.82 x 10-4 2.78 x 10-11
0.206 1.84 x 10-4 3.68 x 10-4 2.49 x 10-11
0.152 1.36 x 10-4 2.72 x 10-4 1.006 x 10-11

2) The experimental procedure for this experiment has you added 5 mL of AgNO3 to 5 mL of 0.0024 M K2CrO4. Is either of these reagents in excess? If so, which one? Elaborate your answer.
1
Expert's answer
2020-12-21T03:38:09-0500



ANSWER1. Ag2​CrO4​⇌2Ag++CrO4​2−
Ksp=(2s)2×s=4s2
Ksp=(1.1×10−12)
S=34
Ksp​=6.5×10−5M
2. AgCl⇌Ag++Cl−
Ksp=S×S ;    Ksp=1.8×10−10
S=s=1.34×10−5M
3. AgBr⇌Ag++Br−
Ksp=S×S ;   Ksp=5×10−13
S=s=0.71×10−6M
4. AgI⇌Ag++I−
Ksp=S×S ;   Ksp=8.3×10−17
S=s=0.9×10−8M
∴ Solubility of Ag2​CrO4​ is highest, so it will precipitate last.